SOHO SCIENCE REPORT No. 8

November 1994

Editors:

Vicente Domingo and Bernhard Fleck

Space Science Department of ESA — ESTEC P.O. Box 299, NL–2200 AG Noordwijk The Netherlands Fax: +31 1719 84698 e-mail: vdomingo@soho.estec.esa.nl bfleck@soho.estec.esa.nl

Contents

1. Planned SOHO Science Working Team (SWT) activities

2. Announced SOHO science related conferences

3. Status of the Project

4. SOHO related bibliography (IV)

5. In Memoriam Philippe Delache (1934-1994)

1. Planned SOHO Science Working Team (SWT) activities

Date  Event  Location Contact

1995

January 23-25 SWT–13  ESTEC  V. Domingo (ESTEC)
April 7-8  VIRGO Science Team meeting Stanford  C. Fröhlich
May 3-5  UVCS Science Team meeting SAO  J. Kohl (SAO)
September 11 Launch party  KSFC  V. Domingo or A. Poland

2. Announced SOHO science related conferences

Date  Event  Location  Contact

1995

April 3-6  4th SOHO Workshop  Pacific Grove, CA  thoeksema@solar.stanford.edu

Helioseismology

April 3-7 EGS XX Hamburg, Germany R.A. Harrison (RAL)
Solar-Terrestrial Physics +  rharrison@solar.stanford.edu
Solar and Heliospheric Physics  B.Foing & R.Marsden (ESTEC)
April 3-7 AGU Chapman Conference: Santa Fe, NM J.Borovsky or R.F.Pfaff
Measurement Techniques for  jborovsky@lanl.gov
Space Plasmas: What Works  u6rfp@lepvax.gsfc.nasa.gov

and What Doesn’t

May 23-27  IAU Colloq. 153  Makuhari, Japan  Takeo Kosugi (NAOJ)
Magnetodynamic Phenomena  (near Tokyo)  tkosugi@solar.stanford.edu
in the Solar Atmosphere  Fax: (+81)-422-34-3742
May 29 - Jun. 2 11th Colloq. on Nagoya, Japan K. Jamashita or T. Watanabe
UV and X-Ray Spectroscopy  yamashita@satio.phys.nagoya-u.ac.jp
of Astrophysical and  watanabe@uvlab.mtk.nao.ac.jp

Laboratory Plasmas

June 25-30  Solar Wind 8  Dana Point Resort, CA DMcComas@solar.stanford.edu
July 9-14  X-Ray and Ultraviolet  San Diego, CA  S. Fineschi
Spectroscopy and Polarimetry II  sfineschi@solar.stanford.edu

(SPIE’s 1995 Int. Symp.)

3. Status of the Project

1. Technical and Programmatic In June 1994 the Flight Model (FM) of the Payload Module (PLM) has undergone successfully its System Functional Test (SFT), with all FMs of the experiments working together. After this test and during July the PLM has concluded its activities at Matra Marconi Space (MMS-UK) in Portsmouth and at the beginning of August was shipped to MMS-F in Toulouse. During August further integration activities (UVCS, CELIAS, PDUs) have taken place. The FM of the Service Module (SVM) also completed its test and received some reworked units and the EQM of the Solid State Recorder, after its specific successful qualification. The intended configuration for Thermal Balance/Thermal Vacuum (TB/TV) test at spacecraft level will be one Solid State Recorder and one Tape Recorder. Mating of the two halves of the FM spacecraft took place for the first time on September 3 and was followed by the first planned and successful System Functional Test (SFT) of the whole FM spacecraft. In the last week of September the spacecraft, again demated, was undergoing further experiment modification activities and preparation for the first (SVM only) TB/TV test. An addendum to the Qualification Review took place in July and highlighted unsolved worries on the functionality of the solar array. A plan of action has been discussed and started with the relevant contractor. The Assembly-Integration-Verification (AIV) activities are now proceeding at MMS-F on double shift and Saturday work to keep the tight schedule leading to an end of April termination of the activities in Europe.

2. Ground Segment The first ground segment compatibility test with NASA’s ground system took place as planned in June with success. The lessons learned and some delays exposed in this exercise are being used to control and plan the ongoing activities in the preparation of the SOHO operations.

3. Launcher In June a Mission Peculiar Critical Design Review (M-CDR) of the SOHO launcher (AC-117) took place at Martin Marietta, San Diego. Martin Marietta successfully launched another Atlas II AS (the same version as SOHO’s) on October 6, 1994. Discussions are ongoing on the revised vibration levels required by the launcher.

4. Experiments The status of recovery actions for UVCS, EIT and CELIAS was the subject of a review by Director of Science (D/SCI) mid-September. UVCS, which suffered some problems in June-July, has recently demonstrated a marked improvement according to the revised recovery schedule. The instrument’s Spectrometer Assembly has been in Alenia-Turin since July and is receiving the new Cross Delay Line (XDL) detectors on schedule. The Mirror Occulter Mechanism is being reworked at GSFC; the FM instrument, with some temporary dummies, will undergo system tests up to December and will then be retro-fitted with all its flight elements. The progress of EIT integration and test activity is excellent. A specific worry on the timely delivery of the flight Solid State Detectors of one of the sensors of CELIAS is being closely monitored, but will not affect the schedule for the TB/TV test (Dec 94). After a status meeting on 15 November it seems that the Solid State Detector production problems are solved. Other instruments have been having some finishing touches to their sensors, elements which showed problems changed, or changes in software are being made. During thermal balance/thermal vacuum tests a fully representative payload will be present, after the latest upgrades and changes in Oct/Nov.

5. Science Operations A first SOHO Science Operations Simulation was successfully held at the Experiment Operations Facility (EOF) at GSFC on 15-17 November. The main purposes were a) to train the scientific groups involved in mission operations and b) to verify the development and readiness status of the overall science operations systems including sending commands, receiving telemetry, planning and implementing mission science. The support of the Experimenter Teams has been excellent and the simulation has proven to be an extremely valuable exercise for each of us (experimenters, Ground System personnel, and Project Office) to get acquainted with the facilities and to detect otherwise unforseeable defects in the planned operations.

6. SOHO Data Archive in Europe Following ESA’s request to the European SOHO science community to submit proposals for building and maintaining a SOHO data archive in Europe, three European institutions have responded:

1. Rutherford Appleton Laboratory, Chilton (UK)

2. Universita di Torino (I)

3. Institut d’Astrophysique Spatiale, Orsay, with CNES, Toulouse (F)

At this stage of the ongoing assessment of the proposals it appears that the three institutions will in principle be able to provide the archive. ESA is discussing with them details of the implementation in order to ensure that the requirements set by the SOHO Science Working Team are fulfilled. It is expected that an implementation plan will be completed by January 1995.

6. Launch date ESA and NASA have mutually agreed to shift the SOHO launch date from the announced July 27 date to 11 September 1995.

4. SOHO related bibliography (IV)

The references given below continue a list of SOHO related publications which was started in SOHO Science Report No. 5. The list is based on the IRS database of ESRIN/ESA as well as on references and pre-/reprints sent to us. If you are aware of publications which you think should appear here, please send us references, abstracts or pre-/reprints. A LaTeX file containing the complete list of SOHO related publications can be obtained on request from bfleck@soho.estec.esa.nl.

Aboudarham, J.: SUMER science database. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 115.

Akin, D.; Horber, R.; Wolfson, J.: Three high duty cycle, space-qualified mechanisms. In: NASA.

Ames Research Center, The 27th Aerospace Mechanisms Symposium. 1994, p. 219. Antiochos, S.K.: Mass flows in coronal loops. In: Mass Supply and Flows in the Solar Corona. Proc.

of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 143.

Antonucci, E.: SOHO contribution to the understanding of mass supply and flows in the solar corona. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 149.

Appourchaux, T.: Effect of space radiations on optical filters. Proceedings of the SPIE, vol.2018, 1993,

p. 80. Bonnet, R.M.: SOHO and the solar community. In: Mass Supply and Flows in the Solar Corona. Proc.

of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 3.

Bürgi, A.; Oetliker, M.; Bochsler, P.; Geiss, J.; Coplan, M.A.: Charge exchange of low energy ions in thin carbon foils. J. Appl. Phys. 68 (6), Sep. 1990, 2547-2554.

Dere, K.: Solar mass flow in fine-scale structures. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 21.

Domingo, V.; Fleck, B.: Activity Report of JOSO WG 6: Coordination of SOHO and ground-based

observations. JOSO Annual Report 1993, p. 100. Domingo, V.; Fleck, B.; Poland, A.I.: The scientific payload of the space-based Solar and Heliospheric

Observatory (SOHO). In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 7.

Esser, R.: Unsolved problems of solar wind expansion: can we learn anything from SOHO? In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 331.

Farquhar, R.W.: Utilization of libration-point orbits, lunar gravity-assist, and Earth-return trajectories. In Israel Society of Aeronautics and Astronautics, 34th Israel AnnualConference on Aerospace Sciences, 1994. p. 238.

Fleck, B.: Cooperation between SOHO and ground-based observatories. JOSO Annual Report 1993,

p. 101. Fleck, B.: The helioseismology instruments on board SOHO: VIRGO, GOLF, MDI. JOSO Annual

Report 1993, p. 103. Fleck, B.; Domingo, V.: SOHO - The Solar and Heliospheric Observatory. In: Solar Magnetic Fields,

eds. M.Schuessler and W.Schmidt, Cambridge University Press, 1994, p. 408. Fleck, B.; Domingo, V.; Poland, A.I.: SOHO science opportunities. In: Solar Coronal Structures, eds.

V.Rusin, P.Heinzel, J.-C.Vial. IAU Colloq. 144, 1994, p. 609. Fleck, B.; Domingo, V.; Poland, A.I.: SOHO operations. In: Solar Coronal Structures, eds. V.Rusin,

P.Heinzel, J.-C.Vial. IAU Colloq. 144, 1994, p. 614. Fleck, B.; Domingo, V.; Poland, A.I.: SOHO: science objectives and capabilities. In: Solar Surface

Magnetism, eds. R.J.Rutten and C.J.Schrijver, NATO ASI Series C 433, Kluwer, 1994, p. 517. Fineschi, S.; Naletto, G.; Nicolosi, P.; Noci, G.; Pernechele, C.; Romoli, M.; Spadaro, D.; Tondello,

G.: Ultraviolet Coronograph Spectrometer (UVCS) for the Solar and Heliospheric (SOHO) Mission. In: International Symposium on Space Optics, Proceedings of the Meeting of the Society of Photo-Optical Instrumentation Engineers, Garmisch D, 1993 (Vol. 2209). (1994)

Fineschi, S.; Romoli, M.: Electron coronal density irregularity from measurements of K-coronal and Lyman lines brightness. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 353.

Fink, D.; Shoup, D.: Application of star identification using pattern matching to space ground systems at GSF. In: NASA. Goddard Space Flight Center, Flight Mechanics/Estimation Theory Symposium, 1994, p. 451.

Goedbloed, J.P.; Poedts, S.; Huysmans, G.T.A.; Halberstadt, G.; Holties, H.; Belien, A.J.C.: Magnetohydrodynamic spectroscopy: Large scale computation of the spectrum of waves in plasmas. Future Generation Computer Systems, vol.10, no.2-3, 1994, p. 339.

Gonin, M.; Kallenbach, R.; Bochsler, P.: Charge exchange of hydrogen atoms in carbon foils at 0.4-120

kev. Rev. Sci. Instrum. 65 (3), March 1994, 648-652. Gonin, M.; Kallenbach, R.; Bochsler, P.: Charge exchange of atoms with high first ionization potentials

in carbon foils at E ≤ 35 kev/u. submitted to NIM B, 1994. Grunes, M.R.; Howard, R.A.; Hoppel, K.; Mango, S.A.; Wang, D.: In: NASA. Goddard Space Flight

Center, The 1994 Space and Earth Science Data Compression Workshop. 1994, p. 11. Habbal, S.R.: Small scale structures in the solar corona. In: Mass Supply and Flows in the Solar

Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 37.

Hartmann, R.: Examination of the response capacity of a silicon pixel detector for ions and molecules

in the energy area 100 to 300 keV. Ph.D. Thesis Tech. Univ. Munich. 1994. Judge, P.G.: The “monochromatic density diagnostic” technique: first detection of multiple density

components in the chromosphere of alpha Tauri. Astrophysical Journal, vol.430, no.1, 1994, p. 351. Kallenbach, R.; Gonin, A.; Bürgi, A.; Bochsler, P.: Charge exchange of hydrogen ions in carbon foils.

Nuclear Instruments and Methods in Physics Research B 83, 1993, 68-72. Kjeldseth-Moe, O.: Working group 1: small scale features. In: Mass Supply and Flows in the Solar

Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 123.

Kjeldseth-Moe, O.; Brynildsen, N.; Brekke, P.; Maltby, P.: Fine structure and the emission filling factor. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 89.

Kjeldseth-Moe, O.; Cheng, C.C.: The extension of explosive events from the transition region to the corona. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 85.

Kopp, R.A.: Working group 3: coronal streamers. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 309.

Mason, H.: Spectroscopic diagnostics for CDS and SUMER. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 111.

Moses, D.; Cook, J.W.: The coronal context of transition region explosive events. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 81.

Naletto, G.; Nicolosi, P.; Pernechele, C.; Poletto, G.; Tondello, G.: VUV Optical Performances of the SOHO Ultraviolet Coronograph Spectrometer. In: Multilayer and Grazing Incidence X-Ray/EUV Optics II. Proceedings of the Meeting of the Society of Photo-Optical Instrumentation Engineers, S. Diego CA, 1993 (Vol. 2011). p. 577-587 (1993).

Neugebauer, M.: Observations of the solar wind from coronal holes. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 319.

Oetliker, M.: Response of a passivated implanted planar silicon (PIPS) detector for heavy ions with energies between 25 an 360 keV. Nuclear Instruments and Methods in Physics Research A 337, 1993, 145-148.

Penn, M.J.; Kuhn, J.R.; Arnaud, J.; Mickey, D.L.; Labonte, B.J.: Coronal electron density measurements using the near-IR Fe XIII emission line. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 185.

Pohjonen, J.: Mounting of SSM’s. Technical Research Centre of Finland, Espoo. 1994. Poland, A.I.; Domingo, V.; Fleck, B.: SOHO operations and ground segment. In: Mass Supply and

Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 13.

Poletto, G.: Understanding solar streamers: the role of SOHO. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 241.

Priest, E.: Working group 2: loops and prominences. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 221.

Reale, F.; Maggio, A.; Ciaravella, A.; Peres, G.: ASAP: a systematic approach to plasma spectral synthesis. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 211.

Romoli, M.; Fineschi, S.: Simulated observations of the electron coronal density irregularity. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 359.

Schultz, C.G.: Coronal streamers’ theory. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 233.

Somov, B.V.: Features of mass supply and flows related with reconnection in the solar corona. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 161.

Spadaro, D.; Leto, P.; Antiochos, S.K.: Observational tests for nonequilibrium ionization in the solar

corona. Astrophysical Journal, vol.427, no.1, 1994, p. 453. Spadaro, D.; Ventura, R.; Martin, R.: UV spectral lines from coronal transients. In: Mass Supply and

Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 365.

Strong, K.T.: Observations of loops and prominences. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 133.

van Ballegooijen, A.A.: Magnetic fine structures in coronal loops. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 31.

Verneta, A.I.; Antonucci, E.; Marocchi, D.: Manifestation of magnetic reconnection in coronal streamer current sheets. In: Mass Supply and Flows in the Solar Corona. Proc. of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 299.

Wu, S.T.: Model development for Ulysses and SOHO. Alabama Univ., Huntsville. Final Report 1993. Wu, S.T.; Song, M.T.; Cheng, C.C.; Dryer, M.: In: Mass Supply and Flows in the Solar Corona. Proc.

of the 2nd SOHO Workshop, eds. B.Fleck, G.Noci, and G.Poletto. Space Science Reviews 70, 1994, p. 167.

In Memoriam

Philippe Delache

1927-1994

October 13 we have been shocked by the unexpected and untimely death of Philippe Delache. He was not merely another CoI interested in the science of VIRGO: indeed, it was his initiative to add the helioseismology to the scientific objectives of the potential ESA space mission DISCO (Dual Irradiance and Solar Constant Observations) during its assessment study in 1981. The objective of DISCO was to measure total and spectral irradiance of the Sun from the far UV to the near IR. Because of the continuity needed for the solar oscillation studies, a halo orbit around L1 was proposed. The helioseismological payload comprised sunphotometers similar to those of VIRGO and a high resolution spectrometer for velocity measurements similar to GOLF. When in December 1982 DISCO was not selected as a mission, the helioseismology instruments, the total solar irradiance and the in situ solarwind measurements were taken over by a new solar mission just starting an assessment study. The Solar High-resolution Observatory (SOHO) for the study of the corona then became the Solar and Heliospheric Observatory. Owing to the much better pointing characteristics of SOHO, it was possible to augment the original DISCO helioseismological payload with a high-resolution instrument. Thus, it is as a consequence of Philippe’s early initiative that a powerful and sophisticated helioseismological payload on the ESA/NASA mission SOHO is now becoming a reality. To honour this and his continuous contributions to the field of helioseismology and solar variability we dedicate VIRGO to the memory of Philippe Delache. Claus Fröhlich and the VIRGO team